# IEEEexample:BSTcontrol

### v2

@@ Abstract
In this paper, we demonstrate that the Ramulator 2.0 simulation results reported in [1] are incorrect and, at the time of the publication of [1], irreproducible. We find that the authors of [1] made multiple trivial human errors in both the configuration and usage of the simulators. We show that by correctly configuring Ramulator 2.0, Ramulator 2.0’s simulated memory system performance actually resembles real system characteristics well, and thus a key claimed contribution of [1] is factually incorrect. We also identify that the DAMOV simulation results in [1] use wrong simulation statistics that are unrelated to the simulated DRAM performance. Moreover, the Mess paper’s artifact repository [4, 5] lacks the necessary sources (simulator code, system configurations, memory traces, etc.) to fully reproduce all the Mess paper’s results. We find that the experiment scripts in [4, 5] use simulator executables and other resources that are neither described in the Mess paper nor found in the artifact repository [4, 5].
Our work corrects [1]’s errors regarding Ramulator 2.0 and identifies important issues in [1]’s memory simulator evaluation methodology. We emphasize the importance of both carefully and rigorously validating simulation results and contacting simulator authors and developers, in true open source spirit, to ensure these simulators are used with correct configurations and as intended. Had we, as developers and maintainers of Ramulator 2.0, been properly contacted before the publication of [1], we could have easily identified the errors in the Mess paper and helped the authors of [1] to avoid publishing factually incorrect results and contributions. We encourage the authors of [1] and the computer architecture community to correct [1]’s errors. This is necessary to prevent the propagation of inaccurate and misleading results, and to maintain the reliability of the scientific record. Our investigation also opens up questions about the integrity of the review and artifact evaluation processes. To aid future work, our source code and scripts are openly available at https://github.com/CMU-SAFARI/ramulator2/tree/mess.
@@ 1 Introduction
Cycle-level DRAM simulators enable research focused on improving the performance, efficiency, and robustness (including security, safety, reliability) of DRAM-based memory systems by providing accurate and flexible models for DRAM and memory controller operations. Ramulator 2.0 [2, 6], (which builds on Ramulator [7, 8]) is a highly modular and extensible cycle-accurate DRAM simulator that enables rapid exploration of new ideas in DRAM-based memory systems.
@@
A MICRO 2024 best paper runner-up publication, A Mess of Memory System Benchmarking, Simulation and Application Profiling [1], which we refer to as “the Mess paper” throughout this paper, with all three artifact badges awarded (including “Reproducible”) proposes a new benchmark to evaluate real and simulated memory system performance. While doing so, it makes some strong negative claims about Ramulator 2.0 and DAMOV [3, 9], which we demonstrate to be incorrect and are due to configuration errors made in the Mess paper.
@@
Fig. 1-(a) and (b) (copied from Fig. 6 and Fig. 4 in [1]) show average memory latency as a function of throughput (used memory bandwidth) for the Mess benchmark using a Ramulator 2.0 simulation and a real system, respectively, as depicted in [1]. [^1] The Mess benchmark consists of two workloads called Stream and Pointer-Chase. These workloads run concurrently. The interval between subsequent memory requests in Stream is controlled using NOPs to vary memory intensity. For example, 1K NOPs between subsequent memory requests lead to a relatively low memory intensity and 0 NOPs between subsequent memory requests lead to a very high memory intensity. Each latency-bandwidth curve in Fig. 1 is constructed by connecting Mess benchmark latency-bandwidth results in increasing memory intensity order such that the leftmost end of the curve corresponds to the smallest memory intensity and the other end of the curve corresponds to the largest memory intensity Mess benchmark configuration. Each curve depicts system performance for a unique mixture of read and write memory requests in Stream ranging from 50% reads and 50% writes to 100% reads and 0% writes. From these two figures, the authors of the Mess paper conclude that Ramulator 2.0 “poorly resembles the actual system performance” (Section I of [1]). They write “The simulated memory latency is unrealistically low and the maximum simulated memory bandwidth is only 126 GB/s which is less than a half of the 292 GB/s measured in the actual system” (Section IV of [1]). They attribute the discrepancy between Mess benchmark results of Ramulator 2.0 simulations and the real system to Ramulator 2.0-induced simulation errors: “This indicates that the main source of the large simulation error is indeed Ramulator 2” (Section 4 of [1]). The authors also show other odd and unexpected results. For example, the DRAM latency results from DAMOV (zsim+Ramulator) [3] (depicted in Section 4 of [1]) are unrealistically low and they seem to remain constant regardless of system load.
@@
In this paper, we show that the Ramulator 2.0 simulation results in the Mess paper are incorrect and, at the time of the publication of [1], irreproducible. Fig. 1-(c) shows our reproduction of the memory latency-bandwidth curves for the Mess benchmark using the open source version of Ramulator 2.0 [2] after correcting the error we found in [1]. [^2] We freely and openly release all sources and data used in this work at https://github.com/CMU-SAFARI/ramulator2/tree/mess. One can reproduce all our graphs using the scripts we provide in that repository. These results clearly show that Ramulator 2.0 does not poorly resemble real system performance contrary to what the Mess paper claims.
@@
We carefully investigate the Mess paper’s open source artifacts [5, 4] and have communicated with the Mess paper’s authors to identify the root causes of the errors in [1]. We identify two major Ramulator 2.0 configuration errors and three other scholarship issues regarding Ramulator 2.0 in the Mess paper. We describe these issues in detail in Sections 2.3 and 2.4. In short, the Ramulator 2.0 results in the Mess paper are subject to gross configuration errors. Therefore, these results do not support the strong negative claims (asserted in the Mess paper) about Ramulator 2.0’s modeling accuracy. [^3] Strong negative assertions made in the Mess paper such as “… the main source of the large simulation error is indeed Ramulator 2…” (Section 4 of [1]) and “… the most complex and trusted memory model shows the highest simulation error…” (Section 4 of [1]) are incorrect because the source of the error is the misconfiguration and misuse of Ramulator 2.0 in the Mess paper. The configuration errors could have been avoided by careful and rigorous validation and/or by contacting Ramulator 2.0 authors and developers (before publishing such incorrect results) to ensure the simulator is correctly used with the correct configuration and as intended.
@@
We also identify that the reason for unrealistically low and seemingly constant DRAM latency results for DAMOV from the Mess paper is that the Mess paper’s authors report simulator statistics that are not updated during the DRAM simulation. Moreover, the Mess paper’s artifact repository [4, 5] does not contain the relevant DRAM simulation statistics from Ramulator for DAMOV. We find that the experiment scripts of the Mess paper use simulator executables and other resources that are neither described in the Mess paper nor found in the Mess paper’s artifact repository [5, 4].
@@
Our work corrects [1]’s errors regarding Ramulator 2.0 and identifies important scholarship issues in [1]’s memory simulator evaluation methodology. We emphasize the importance of contacting simulator authors and developers (especially before publishing), in true open source spirit, to ensure these simulators are used with correct configurations and as intended. We encourage the authors of the Mess paper and the computer architecture community to correct [1]’s errors to prevent propagation of inaccurate and misleading results and to maintain the reliability of the scientific record. Our investigation also opens up questions about the integrity of the review and artifact evaluation processes.
@@ 2.1 The Mess Benchmark
The Mess benchmark aims to characterize a memory system based on a set of bandwidth-latency curves. To this end, the Mess paper introduces the Mess benchmark. The Mess benchmark is made up of two concurrently running workloads, Stream and Pointer Chase [10, 11]. Each curve is given by running these workloads with a pre-configured fraction of read and write requests for the Stream, ranging from 50% reads and 50% writes to 100% reads. Each point on a curve is given by a run of the benchmark with varying memory intensity. This is achieved by varying the interval between subsequent Stream requests by inserting NOPs. For example, 1K NOPs between subsequent memory requests lead to a relatively low memory intensity, and 0 NOPs between subsequent memory requests lead to a very high memory intensity.
@@
Fig. 2 (Fig. 1-a plotted again for the reader’s convenience) shows the bandwidth-latency curves presented in [1] (Fig. 6 in [1]) for Ramulator 2.0 obtained with the Mess benchmark. This result unexpectedly and oddly shows that the simulated memory system 1) uses significantly smaller bandwidth than what is available (depicted as the “Max. theoretical BW” in the figure) and 2) yields a relatively small memory access latency.
Fig. 2: Latency-bandwidth curves for Ramulator 2.0 (copied from Fig. 6 in [1])
We note that it was impossible to reproduce any Ramulator 2.0 result from the Mess paper artifacts available at the time of the publication and artifact evaluation of the Mess paper [5, 4]. Thus, we contacted the authors of [1] on 16 October 2024 to obtain 1) the Ramulator 2.0 source code and 2) the Mess benchmark memory traces used in Ramulator 2.0 simulations reported in [1]. The missing sources needed to replicate the Mess paper’s Ramulator 2.0 results were eventually uploaded to the Mess Github repository [5] following our contact on 18 July 2025, 10 months after the Mess paper was awarded (on 8 September 2024) all three artifact evaluation badges.
@@ 2.3 Fundamental Configuration Errors in [1]
We identify two fundamental configuration errors in the Mess paper’s Ramulator 2.0 simulations. We attribute the discrepancy in bandwidth and memory access latency between the Mess paper’s Ramulator 2.0 simulations and real system evaluation to these fundamental errors. We also identify three other scholarship issues.
First, the Mess paper’s scripts (in the newly-uploaded source code on 18 July 2025) instantiate and use a single-channel DDR5 Ramulator 2.0 memory system. The authors extrapolate single-channel DDR5 results to 16-channel DDR5 results by multiplying the observed bandwidth by 8 instead of 16. This is because they incorrectly assume that a DDR5 channel in Ramulator 2.0 is 64 bits wide, consisting of two independent 32-bit channels. However, the source code of Ramulator 2.0 clearly shows that a DDR5 channel is 32 bits wide [^4] The m_channel_width parameter in DDR5.cpp in the Ramulator 2.0 repository [6]. in Ramulator 2.0. The Mess paper misleadingly presents (extrapolated) 8-channel (32-bit each) DDR5 results as 16-channel (32-bit each) DDR5 results, leading to incorrect claims like “…the maximum simulated memory bandwidth is only 126 GB/s which is less than a half of the 292 GB/s measured in the actual system” in Section 4 of [1]. The maximum bandwidth of an 8-channel DDR5-4800 [12] is 153.6 GB/s [12], not 307 GB/s. Thus, a bandwidth of more than 153.6 GB/s should not be expected of an 8-channel DDR5-4800 configuration. We show that a correctly configured 16-channel DDR5-4800 Ramulator 2.0 simulation achieves 281.1 GB/s bandwidth (as depicted in Fig. 1-c; see Section 4 for a detailed description).
Second, the Mess paper’s scripts use the SimpleO3 frontend of Ramulator 2.0 [6] with unrealistically low latency configurations compared to a real system. For example, they configure 1) the cache latency to be 0 CPU clock cycles and 2) each CPU core to have 1024 miss-status holding registers (MSHRs). Compared to a real CPU, such a configuration has 1) significantly smaller round-trip latency between a core and DRAM, 2) significantly lower queuing delay for memory requests when memory intensity is high. Thus, the Mess paper’s claim about Ramulator 2.0 that “…the simulated memory latency is unrealistically low…” (Section 4 of [1]) is unfair since the Mess paper’s unrealistically low-latency configuration of Ramulator 2.0 is in part responsible for such low latency.
@@ 2.4 Scholarship Issues in [1]
First, the Mess paper’s authors did not contact us regarding the results they obtained with their configuration of Ramulator 2.0 before the Mess paper was published. Therefore, we could not help them correct the issues before it was in scientific record. Their misconfiguration of the number of channels in DDR5 could have been easily resolved by emails and collegial discussions which are in the spirit of open-source development.
Second, the Mess paper does not disclose the unrealistically low latency configurations of the SimpleO3 frontend of Ramulator 2.0 in the paper. The Ramulator 2.0 configuration file available in the artifact repository also does not include any of these configurations. After the Mess paper’s authors uploaded their Ramulator 2.0 source code (10 months after the Mess paper was awarded all three artifact evaluation badges), we found (via email exchanges with the Mess paper’s authors) that they implement the unrealistically low latency configurations by directly modifying the source code of Ramulator 2.0, without any changes in the Ramulator 2.0 configuration file.
Third, the Mess paper uses different Mess benchmark methodologies for real systems and simulators without disclosing this difference in the paper. [^5] Identified via email exchanges with the Mess paper’s authors. For real systems (e.g., Fig. 5-a in [1]), the Mess benchmark is run as described in the Mess paper (i.e., by concurrently running pointer-chasing and Stream workloads and measuring the memory latency of pointer-chasing workload’s memory accesses). For simulators (e.g., Figures 5-b, -c, -d, -e, and -f in [1]), the authors do not run a pointer-chasing workload and they measure the average memory latency for only the Stream workload. Thus, the Mess paper’s authors compare apples to oranges by running and measuring different workloads in simulators and in real systems. And, this difference in simulation and real system evaluation methodologies is not disclosed in the Mess paper.
Through our comprehensive investigation of [1]’s Ramulator 2.0 artifacts, we conclude that the Ramulator 2.0 results in [1] are subject to multiple serious configuration and methodological errors that could have been easily avoided by careful and rigorous validation and/or by properly contacting the authors of Ramulator 2.0 in the true open-source spirit before the submission (and certainly before publication) of the Mess paper.

### v3

# Cleaning up the Mess: Re-Evaluating the Real-System Modeling Accuracy of Ramulator 2.0

@@ Abstract
In this paper, we show that the Mess paper [2] has 1) demonstrable technical misconfigurations, 2) methodological errors in interpreting simulation statistics, and 3) an incomplete artifact that makes its key results irreproducible. We demonstrate that the Ramulator 2.0 simulation results reported in [2] are incorrect due to multiple configuration errors instead of inherent simulation inaccuracy claimed by the Mess paper. We show that by correctly configuring Ramulator 2.0, Ramulator 2.0’s simulated memory system performance actually resembles real system characteristics well, and thus a key claimed contribution of [2] is factually incorrect. We also identify that the DAMOV simulation results in [2] use wrong simulation statistics that are unrelated to the simulated DRAM performance. By using correct DRAM simulation statistics, we show that DAMOV’s simulated DRAM latency is not constant, in contrast to [2]’s claim. Moreover, the Mess paper’s artifact repository [3, 17] lacks the necessary sources (simulator code, system configurations, memory traces, etc.) to fully reproduce all the Mess paper’s results. We find that the experiment scripts in [3, 17] use simulator executables and other resources that are neither described in the Mess paper nor found in the artifact repository [3, 17].
Our work identifies important issues in [2]’s memory simulator evaluation methodology regarding Ramulator 2.0 and DAMOV. We present results that validate the real-system modeling accuracy of Ramulator 2.0, and describe the reasons why [2]’s results with respect to these two simulators are incorrect. We emphasize the importance of carefully and rigorously validating simulation results to avoid publishing factually incorrect results and contributions. We strongly encourage the computer architecture community to consider our corrections to the Ramulator 2.0 and DAMOV results of the Mess paper [2] to prevent the propagation of inaccurate and misleading results and to maintain the reliability of the scientific record. Our investigation also aims to stimulate discussion on artifact evaluation practices and on mechanisms for correcting results and artifacts after publication. To aid future works and reproduction of all our results, we open source all our code and scripts. We also discuss best practices and add sanity checks to aid users and developers of simulation tools.
11footnotetext: F. Nisa Bostancı, Haocong Luo, and Ataberk Olgun are co-primary authors.
@@ 1 Introduction
Cycle-level DRAM simulators enable research focused on improving the performance, efficiency, and robustness (including security, safety, reliability) of DRAM-based memory systems by providing accurate and flexible models for DRAM and memory controller operations. Ramulator 2.0 [16, 7] (which builds on Ramulator [13, 8]), is a highly modular and extensible cycle-accurate DRAM simulator that enables rapid exploration of new ideas in DRAM-based memory systems. DAMOV [30, 5] is a simulation framework that integrates Ramulator [13, 8] into ZSim CPU simulator [32] to enable fast, scalable, and cycle-level system simulation.
@@
A MICRO 2024 best paper runner-up publication, A Mess of Memory System Benchmarking, Simulation and Application Profiling [2], which we refer to as “the Mess paper” throughout this paper, with all three artifact badges awarded (including “Reproducible”) proposes a new benchmark to evaluate real and simulated memory system performance. While doing so, it makes some strong negative claims about Ramulator 2.0 and DAMOV, which we demonstrate to be incorrect and are due to configuration and simulator usage errors made in the Mess paper.
@@
Fig. 1-(a) and (b) (copied from Fig. 6 and Fig. 4 in [2]) show average memory latency as a function of throughput (used memory bandwidth) for the Mess benchmark using a Ramulator 2.0 simulation and a real system, respectively, as depicted in [2]. [^1] The Mess benchmark consists of two workloads called Stream and Pointer-Chase. These workloads run concurrently. The interval between subsequent memory requests in Stream is controlled using NOPs to vary memory intensity. For example, 1K NOPs between subsequent memory requests lead to a relatively low memory intensity and 0 NOPs between subsequent memory requests lead to a very high memory intensity. Each bandwidth–latency curve in Fig. 1 is constructed by connecting Mess benchmark bandwidth–latency results in increasing memory intensity order such that the leftmost end of the curve corresponds to the smallest memory intensity and the other end of the curve corresponds to the largest memory intensity Mess benchmark configuration. Each curve depicts system performance for a unique mixture of read and write memory requests in Stream ranging from 50% reads and 50% writes to 100% reads and 0% writes. From these two figures, the authors of the Mess paper conclude that Ramulator 2.0 “poorly resembles the actual system performance” (Section I of [2]). They write “The simulated memory latency is unrealistically low and the maximum simulated memory bandwidth is only 126 GB/s which is less than a half of the 292 GB/s measured in the actual system” (Section IV of [2]). They attribute the discrepancy between Mess benchmark results of Ramulator 2.0 simulations and the real system to Ramulator 2.0-induced simulation errors: “This indicates that the main source of the large simulation error is indeed Ramulator 2” (Section 4 of [2]). The authors also show other unexpected results. For example, the DRAM latency results from DAMOV (ZSim+Ramulator) [5] (depicted in Section 4 of [2]) are unrealistically low and they seem to remain constant regardless of system load.
@@
In this paper, we show that the Ramulator 2.0 simulation results in the Mess paper are incorrect and, at the time of the publication of [2], irreproducible. Fig. 1-(c) shows our reproduction of the memory bandwidth–latency curves for the Mess benchmark using the open source version of Ramulator 2.0 [16] after correcting the error we found in [2]. These results clearly show that Ramulator 2.0 does not poorly resemble real system performance contrary to what the Mess paper claims.
@@
We carefully investigate the Mess paper’s open source artifacts [17, 3] and have communicated with the Mess paper’s authors to identify the root causes of the errors in [2]. We identify two major Ramulator 2.0 configuration errors and two other scholarship issues regarding Ramulator 2.0 in the Mess paper. We describe these issues in detail in Sections 2.4 and 2.5. In short, the Ramulator 2.0 results in the Mess paper are subject to configuration errors. Therefore, these results do not support the negative claims (asserted in the Mess paper) about Ramulator 2.0’s modeling accuracy. [^2] Strong negative assertions made in the Mess paper such as “… the most complex and trusted memory model shows the highest simulation error…” (Section 4 of [2]) are incorrect because the source of the error is the misconfiguration and misuse of Ramulator 2.0 in the Mess paper. Ramulator 2.0 can be correctly configured by careful and rigorous validation and/or by contacting Ramulator 2.0 authors and developers after observing unexpected results, to ensure that the simulator is used correctly and as intended. [^3] The Mess paper’s authors verified via email exchanges that they intentionally did not contact the developers of the memory system simulators evaluated in the Mess paper before publication of the paper.
@@
We also identify that the reason for unrealistically low and seemingly constant DRAM latency results for DAMOV from the Mess paper is that the Mess paper’s authors report simulator statistics that are not updated during the DRAM simulation. Moreover, the Mess paper’s artifact repository [3, 17] does not contain the relevant DRAM simulation statistics from Ramulator for DAMOV. We find that the experiment scripts of the Mess paper use simulator executables and other resources that are neither described in the Mess paper nor found in the Mess paper’s artifact repository [17, 3]. We independently evaluate DAMOV’s simulated memory system performance using the open source version of DAMOV [5]. We show that, when using the correct DRAM simulation statistics from Ramulator, 1) the simulated DRAM latency does not remain constant in contrast to the results claimed by the Mess paper [2], and 2) DAMOV’s bandwidth–latency curves resemble real system characteristics.
@@
Our work provides corrections to the Ramulator 2.0 and DAMOV results of the Mess paper [2] and identifies important scholarship issues in the memory simulator evaluation methodology of the Mess paper [2]. We emphasize the importance of contacting simulator authors and developers (especially before publishing), in true open source spirit, to ensure these simulators are used with correct configurations and as intended. We strongly encourage the computer architecture community to consider our corrections to the Ramulator 2.0 and DAMOV results of the Mess paper [2] to prevent the propagation of inaccurate and misleading results and to maintain the reliability of the scientific record. Our investigation also aims to stimulate discussion on artifact evaluation practices and on mechanisms for correcting results and artifacts after publication. We discuss four best practices and add sanity checks to aid users and developers of simulation tools (§6). To aid future works and reproduction of all our results, we open source all our code and scripts at https://github.com/CMU-SAFARI/Cleaning-up-the-Mess.
@@ 2.1 The Mess Benchmark
The Mess benchmark aims to characterize a memory system based on a set of bandwidth–latency curves. To this end, the Mess paper introduces the Mess benchmark. The Mess benchmark is made up of two concurrently running workloads, Stream and Pointer Chase [35, 34]. Each curve is given by running these workloads with a pre-configured fraction of read and write requests for the Stream, ranging from 50% reads and 50% writes to 100% reads and 0% writes. Each point on a curve is given by a run of the benchmark with varying memory intensity. This is achieved by varying the interval between subsequent Stream requests by inserting NOPs. For example, 1K NOPs between subsequent memory requests lead to a relatively low memory intensity, and 0 NOPs between subsequent memory requests lead to a very high memory intensity.
@@
Fig. 2 (Fig. 1-a plotted again for the reader’s convenience) shows the bandwidth–latency curves presented in [2] (Fig. 6 in [2]) for Ramulator 2.0 obtained with the Mess benchmark. This result unexpectedly and oddly shows that the simulated memory system 1) uses significantly smaller bandwidth than what is available (depicted as the “Max. theoretical BW” in the figure) and 2) yields a relatively small memory access latency.
Fig. 2: Bandwidth–latency curves for Ramulator 2.0 (copied from Fig. 6 in [2])
2.3 Reproducibility Issues in [2]
We note that it was not possible to reproduce any Ramulator 2.0 result from the Mess paper artifacts available at the time of the publication and artifact evaluation of the Mess paper [17, 3]. Thus, we contacted the authors of [2] on 16 October 2024 to obtain 1) the Ramulator 2.0 source code and 2) the Mess benchmark memory traces used in Ramulator 2.0 simulations reported in [2]. The missing sources needed to replicate the Mess paper’s Ramulator 2.0 results were eventually uploaded to the Mess GitHub repository [17] following our contact on 18 July 2025, 10 months after the Mess paper was awarded (on 8 September 2024) all three artifact evaluation badges.
@@ 2.4 Fundamental Configuration Errors in [2]
We identify two fundamental configuration errors in the Mess paper’s Ramulator 2.0 simulations. We attribute the discrepancy in bandwidth and memory access latency between the Mess paper’s Ramulator 2.0 simulations and real system evaluation to these fundamental errors. We also identify two other scholarship issues.
First, the Mess paper’s scripts (in the newly-uploaded source code on 18 July 2025) instantiate and use a single-channel DDR5 Ramulator 2.0 memory system. The authors extrapolate single-channel DDR5 bandwidth measurements to 16-channel DDR5 bandwidth measurements by multiplying the observed bandwidth by 8 instead of 16. This is because they incorrectly assume that a DDR5 channel in Ramulator 2.0 is 64 bits wide, consisting of two independent 32-bit channels. However, the source code of Ramulator 2.0 clearly shows that a DDR5 channel is 32 bits wide [^4] The m_channel_width parameter in DDR5.cpp in the Ramulator 2.0 repository [7]. in Ramulator 2.0, which is consistent with JEDEC DDR5 standards [10, 12, 11]. [^5] The JEDEC DDR5 SDRAM component [10] and UDIMM module [12] standards consistently use the term “channel”. The JEDEC DDR5 RDIMM standard [11] uses the terms “channel” and “sub-channel” interchangeably. However, both “channel” and “sub-channel” in these JEDEC DDR5 documents [10, 12, 11] are 32-bit wide. The Mess paper misleadingly presents (extrapolated) 8-channel (32-bit each) DDR5 results as 16-channel (32-bit each) DDR5 results, leading to incorrect claims like “…the maximum simulated memory bandwidth is only 126 GB/s which is less than a half of the 292 GB/s measured in the actual system” in Section 4 of [2]. The maximum bandwidth of an 8-channel DDR5-4800 [10] is 153.6 GB/s [10], not 307 GB/s. Thus, a bandwidth of more than 153.6 GB/s should not be expected of an 8-channel DDR5-4800 configuration. We show that a correctly configured 16-channel DDR5-4800 Ramulator 2.0 simulation achieves 281.1 GB/s bandwidth (as depicted in Fig. 1-c; see Section 4 for a detailed description).
Second, the Mess paper’s scripts use the SimpleO3 frontend of Ramulator 2.0 [7] with unrealistically low latency configurations compared to a real system. For example, they configure 1) the cache latency to be 0 CPU clock cycles and 2) each CPU core to have 1024 miss-status holding registers (MSHRs). Compared to a real CPU, such a configuration has 1) significantly smaller round-trip latency between a core and DRAM, and 2) significantly lower queuing delay for memory requests when memory intensity is high. About Ramulator 2.0’s simulation accuracy, the Mess paper claims that “…the simulated memory latency is unrealistically low…” (Section 4 of [2]). Because the unrealistic cache setup is in part responsible for the low latency observed in [2], the cache setup in [2] constitutes a fundamental configuration error. Thus, the Ramulator 2.0 simulation accuracy claim made in the Mess paper [2] is incorrect.
@@ 2.5 Scholarship Issues in [2]
First, the Mess paper does not disclose the unrealistically low latency configurations of the SimpleO3 frontend of Ramulator 2.0 in the paper. The Ramulator 2.0 configuration file available in the artifact repository also does not include any of these configurations. After the Mess paper’s authors uploaded their Ramulator 2.0 source code, we found (via email exchanges with the Mess paper’s authors) that they implement the unrealistically low latency configurations by directly modifying the source code of Ramulator 2.0, without any documented changes.
Second, the Mess paper uses different Mess benchmark methodologies for real systems and simulators without disclosing this difference in the paper. [^6] Identified via email exchanges with the Mess paper’s authors. For real systems (e.g., Fig. 5-a in [2]), the Mess benchmark is run as described in the Mess paper (i.e., by concurrently running pointer-chasing and Stream workloads and measuring the memory latency of the pointer-chasing workload’s memory accesses). For simulators (e.g., Figures 5-b, -c, -d, -e, and -f in [2]), the authors do not run a pointer-chasing workload and they measure the average memory latency for only the Stream workload. Thus, the Mess paper’s authors compare apples to oranges by running and measuring different workloads in simulators and in real systems. And, this difference in simulation and real system evaluation methodologies is not disclosed in the Mess paper.
Through our comprehensive investigation of [2]’s Ramulator 2.0 artifacts, we conclude that the Ramulator 2.0 results in [2] are subject to multiple configuration and methodological errors that could have been avoided by careful and rigorous validation.